Find values for m and b in the following system so that the solution to the system is (-6, 8).
5x + 7y = b
mx + y = 22
m = , b =

Answer by scott8148(6628)

substituting __ 5(-6)+7(8)=b __ -30+56=b

substituting __ m(-6)+(8)=22 __ -6m=14